eviant:

Let f(x) = -2x - 7 and g(x) = -4x + 6. Find (f*g)(-5)

eviant:

@mhchen

mhchen:

I think (f*g)(-5) = f(-5)*g(-5). so first find f(-5)

eviant:

10x+35

mhchen:

uh well f(-5) = -2(-5) - 7 = 10 - 7 = 3

mhchen:

I'm not sure how you got 10x + 35, Did you see what I did? Try to find g(-5) now,

eviant:

-4(-5)+6 20+6 26?

mhchen:

Yup so all you have to do now is multiply f(-5) and g(-5) together 3 * 26

eviant:

78?

mhchen:

ye

eviant:

3 –6 26 –59

mhchen:

oh are those the options?

eviant:

^^those are the choices

mhchen:

HMMMmmmmm

mhchen:

okay I'm gonna assume (f*g)(x) is (fog)(x) or f(g(x)) a composite function

mhchen:

basically first you find g(-5) which you already did. It was 26 Then you plug 26 into f(x)

mhchen:

Can you find f(26)?

eviant:

@mhchen -2(26)-7=-59?

mhchen:

yep